# AMBER: Aerial deployable gripping crawler with compliant microspine for canopy manipulation

### v1

@@ Abstract
This paper presents an aerially deployable crawler designed for adaptive locomotion and manipulation within tree canopies. The system combines compliant microspine-based tracks, a dual-track rotary gripper, and an elastic tail, enabling secure attachment and stable traversal across branches of varying curvature and inclination. Experiments demonstrate reliable gripping up to 90° body roll and inclination, while effective climbing on branches inclined up to 67.5°, achieving a maximum speed of 0.55 body lengths per second on horizontal branches. The compliant tracks allow yaw steering of up to 10°, enhancing maneuverability on irregular surfaces. Power measurements show efficient operation with a dimensionless cost of transport over an order of magnitude lower than typical hovering power consumption in aerial robots. Integrated within a drone–tether deployment system, the crawler provides a robust, low-power platform for environmental sampling and in-canopy sensing, bridging the gap between aerial and surface-based ecological robotics.
@@ I Introduction
Figure 1: Mission concept of the AMBER framework: the drone deploys the crawler onto a branch, perches, and the crawler traverses the branch to perform sampling before being retrieved by the drone.
@@
To address these limitations, we present AMBER, a drone-deployed crawler for branch traversal and direct sampling. It features a dual-track mechanism with compliant microspines and a spring–damping suspension, enabling adaptive gripping on natural surfaces. Deployed via an aerial tether system, it combines energy-efficient perching with the mobility needed for in-canopy exploration and sensing, including payload transport along branches of varying diameters and inclinations and transitions onto trunks.
@@
Existing climbing robots such as RiSE [20] and TreeBot [21] demonstrate impressive vertical mobility using gecko-like adhesion and microspine grasping, yet they are either too heavy for aerial deployment or do not demonstrate complete crawling or gripping around the branch, which is useful for sampling epiphytes. Continuous-motion systems using wheels or tracks [22, 23, 24], or trunk-enclosing designs [25], remain confined to smooth or vertical surfaces. To overcome these limitations, AMBER introduces a novel framework for tree branch and trunk locomotion and manipulation, enabling rapid, aerial-deployable access for environmental sampling in complex canopy environments. Aerial deployment enables rapid positioning of the crawler within the canopy, while the tether allows the drone to perch securely on a branch and provides safe recovery for the crawler, ensuring continued operation after a fall. With a sampling mechanism, the crawler can traverse the tree structure and collect biological or environmental samples, as shown in Fig. 2. After sampling, the crawler is retrieved through the tether, providing a safe and cost-effective approach for in-canopy exploration and analysis. This capability also supports targeted ecological monitoring, such as detecting or limiting the spread of invasive pests and diseases, by positioning sensors on specific parts of the canopy.
@@
  • Aerial-deployable framework: A novel aerial–terrestrial system that enables safe, rapid, and autonomous deployment of a crawler into the canopy via a tether. This framework allows recovery in case of failure and localized, cost-effective sampling of plants and epiphytes, paving the way for scalable, non-invasive canopy studies.
@@ II-C Compliant microspines
Figure 3: Track body (in grey) illustrating the actuation of the spine carriers. (a) Intermediate chain link, (b) Retainer. (c) Compliant microspine design consisting of a frame, spring, and spine carrier, allowing lateral movement that facilitates crawler maneuverability.
@@ II-E Wheeled tail
Tails are an evolutionary feature that assist animals with balance and climbing (cite biological examples). Robotic studies have shown that adding a tail can significantly enhance locomotion performance [27]. Building on this insight, we hypothesize that incorporating a tail into AMBER could improve branch-climbing ability by increasing grip and overall stability.
The tail, designed to be 100 mm long, includes a wheel with a tunable spring–damping system, which reduces friction and improves stability during climbing. This design allows the wheel to apply the optimal pushing force, enabling secure, adaptive gripping on branches and enhancing overall climbing performance. Inspired by prior work on climbing robots such as Tbot [23] and soft-legged systems [28], we build on these locomotion principles while integrating the wheel specifically into the tail.
@@ III Experimental Results
The robot and its features were validated through a series of comprehensive gripping and crawling tests, which were designed to evaluate and characterize the system’s complete operational performance.
@@ III-A Gripping
For these tests, a branch with a diameter of 95 mm was used. The gripping angle was evaluated prior to testing the performance at various inclinations, yaw, and roll in order to maximize the normal force. The angle affects both the proximity of the crawler to the branch and the orientation at which the needles engage the branch. For this specific branch, the gripping angle was set to 84.4°. A potential method to automatically adjust the angle based on branch characteristics would be to use a force sensor to detect when sufficient grip is achieved by the needles.
@@
As shown in Fig. 5(a), maximum pull-off forces were recorded for configurations with and without the tail, confirming the tail’s importance on steep branches. Pull-off forces decreased with increasing pitch due to the weight-induced pitching moment. This is consistent with observations in arboreal quadrupeds, where tails act as stabilizing appendages that improve balance and mobility on narrow or inclined substrates [29]. Without the tail, the robot’s gripping performance was limited to branches with a maximum steepness of 45°. The elastically loaded tail redistributed the load, increased gripping forces, and enabled stable attachment on branches up to 90°, supporting payloads of up to 10 N.
@@ III-C Power consumption & Cost of Transport
Tables I and II summarize the results. No-load corresponds to the robot running without payload, with an average power of 9.85 W. Perching represents the robot holding position without movement, consuming only 3.15 W, nearly zero compared to continuous motion. Crawling on a horizontal branch required 13.35 W, while peak power under maximum load reached 34.2 W. The COT during horizontal crawling was 23.4, increasing to 64.6 at a 67.5° incline, reflecting the higher energetic cost of climbing steeper surfaces.
@@
For comparison, UAV flight requires continuous thrust and correspondingly higher power. For example, a DJI F450 drone consumes roughly 300 W while hovering with a 0.933 kg payload [16], whereas the crawler consumes only 3.15 W when perched and 13.35 W during horizontal crawling. These results demonstrate that the crawler achieves more than one order of magnitude lower power expenditure than typical aerial hovering, highlighting its efficiency for slow, contact-based locomotion.
IV Conclusions
This work presented an aerially deployable crawler designed for adaptive locomotion and manipulation within tree canopies. The system combines compliant microspine-based tracks, a dual-track rotary grasper, and an elastic tail, enabling secure attachment and traversal on branches of varying diameters and inclinations.
Experiments demonstrated stable gripping on rolls up to 90° and climbing on branches up to 67.5° inclination, with a 33% success rate. On horizontal branches, the robot achieved an average speed of 8.33 cm/s (0.55 body lengths/s), which is comparable to or higher than state-of-the-art crawlers. The compliant microspine mechanism enabled yaw steering up to 10°, enhancing maneuverability on irregular surfaces, while the elastic tail improved stability during steep ascents and perching.
Power measurements indicated low energy requirements across all modes: 3.15 W while perched, 13.35 W during horizontal crawling, and up to 18.46 W on steep inclines. The corresponding Cost of Transport ranged from 23.4 on horizontal branches to 64.6 at 67.5°, demonstrating efficient operation even on challenging terrain, with power consumption over an order of magnitude lower than that of a hovering drone.
Deployed through a drone–tether system, the crawler can be safely positioned, anchored, and retrieved, enabling in-canopy exploration, environmental sensing, and sample collection. Future work will focus on enhancing tail-assisted stability through force- or geometry-aware control of the gripping angle. Additionally, lightweight sensing and sampling tools will be integrated to extend autonomous operation within forest canopies, enabling drone implementation and full mission execution in real-world field deployments.

### v2

# AMBER: A tether-deployable gripping crawler with compliant microspines for canopy manipulation

@@ Abstract
This paper presents an aerially deployable crawler designed for adaptive locomotion and manipulation within tree canopies. The system combines compliant microspine-based tracks, a dual-track rotary gripper, and an elastic tail, enabling secure attachment and stable traversal across branches of varying curvature and inclination. Experiments demonstrate reliable gripping up to 90° body roll and inclination, while effective climbing on branches inclined up to 67.5°, achieving a maximum speed of 0.55 body lengths per second on horizontal branches. The compliant tracks allow yaw steering of up to 10°, enhancing maneuverability on irregular surfaces. Power measurements show efficient operation with a dimensionless cost of transport over an order of magnitude lower than typical hovering power consumption in aerial robots. The crawler provides a robust, low-power platform for environmental sampling and in-canopy sensing. The aerial deployment is demonstrated at a conceptual and feasibility level, while full drone–crawler integration is left as future work.
@@ I Introduction
Figure 1: Conceptual mission illustration (not to scale). A drone deploys a tethered crawler onto a branch. Full aerial–crawler integration is future work.
@@
To address these limitations, we present AMBER, a tether-deployed crawler for branch traversal and direct sampling. It features a dual-track mechanism with compliant microspines and a spring–damping suspension, enabling adaptive gripping on natural surfaces. As shown in Fig. 1, it can reach the canopy via aerial tether deployment. It combines energy-efficient perching with the mobility needed for in-canopy exploration and sensing, including payload transport along branches of varying diameters and inclinations and transitions onto trunks.
@@
Existing climbing robots such as RiSE [26] and TreeBot [11] demonstrate impressive vertical mobility using gecko-like adhesion and microspine grasping, yet they are either too heavy for aerial deployment or do not demonstrate complete crawling or gripping around the branch, which is useful for sampling epiphytes. Continuous-motion systems using wheels or tracks [2, 15, 14], or trunk-enclosing designs [8], remain confined to smooth or vertical surfaces. To overcome these limitations, AMBER introduces a novel framework for tree branch and trunk locomotion and manipulation, enabling rapid, aerial-deployable access for environmental sampling in complex canopy environments. Aerial deployment enables rapid positioning of the crawler within the canopy, while the tether allows the drone to perch securely on a branch and provides safe recovery for the crawler, ensuring continued operation after a fall. The crawler is designed to accommodate lightweight sampling modules; however, this work focuses on locomotion and gripping performance. After sampling, the crawler is retrieved through the tether, providing a safe and cost-effective approach for in-canopy exploration and analysis. This capability also supports targeted ecological monitoring, such as detecting or limiting the spread of invasive pests and diseases, by positioning sensors on specific parts of the canopy.
@@
  • Tether-deployable canopy locomotion framework: A lightweight crawler designed for deployment and recovery using an aerial tether, enabling safe positioning and retrieval in tree canopies.
@@ II-C Compliant microspines
Figure 3: Track body components for spine carrier actuation: (a) intermediate chain link, (b) retainer closing the chain insertion slot during assembly, (c) spine carrier, and (d) track body.
Figure 4: Compliant microspine design consisting of a frame, spring, and spine carrier, allowing lateral movement that facilitates crawler maneuverability.
@@ II-E Wheeled tail
Long and mobile tails in arboreal animals improve stability when moving on narrow or precarious substrates by modulating body angular momentum and promoting balance during locomotion [29]. Robotic studies have shown that adding a tail can significantly enhance locomotion performance [25]. Building on this insight, we hypothesize that incorporating a tail into AMBER could improve branch-climbing ability by increasing grip and overall stability.
To minimize mass and control complexity, a passive tail-mounted wheel was selected, prioritizing robustness over optimal force generation. The tail, approximately 100 mm long, integrates a wheel with a tunable spring–damping system that reduces friction and improves stability during climbing. This compliance allows the wheel to passively regulate the applied pushing force, enabling secure and adaptive branch gripping and enhancing overall climbing performance. Inspired by prior climbing robots such as Tbot [15] and soft-legged systems [24], this design adapts established locomotion principles to a wheel-based tail architecture..
@@ III Experimental Results
The robot and its features were validated through a series of comprehensive gripping and crawling tests designed to characterize the system’s complete operational performance across the three principal orientation axes. In this context, pitch refers to the inclination of the branch relative to gravity, roll denotes the robot’s rotation around the branch circumference, and yaw describes the steering misalignment between the crawler’s longitudinal axis and the branch.
@@ III-A Gripping
For these tests, a branch with a diameter of 95 mm was used. The gripping angle was evaluated prior to testing the performance at various inclinations, yaw, and roll in order to maximize the normal force. The angle affects both the proximity of the crawler to the branch and the orientation at which the needles engage the branch. For this specific branch, the maximum pull-off force was recorded at 84.4°, as shown in Fig. 5, and this angle was therefore used for subsequent tests. A potential method to automatically adjust the angle based on branch characteristics would be to use a force sensor to detect when sufficient grip is achieved by the needles.
@@
As shown in Fig. 6(a), maximum pull-off forces were recorded for configurations with and without the tail, confirming the tail’s importance on steep branches. Pull-off forces decreased with increasing pitch due to the weight-induced pitching moment. This is consistent with observations in arboreal quadrupeds, where tails act as stabilizing appendages that improve balance and mobility on narrow or inclined substrates [28]. The elastically loaded tail redistributed the load, increasing the maximum normal pull-off force by up to 15% on inclinations below 45° (Fig. 6(a)) and enabling stable attachment on branches between 45° and 90°, which was not possible without the tail, while supporting payloads of up to 10 N. The wheel relies on friction rather than grasping and is therefore sensitive to surface roughness and curvature.
@@ III-C Power consumption & Cost of Transport
Tables I and II summarize the results. No-load corresponds to the robot running without payload, with an average power of 9.85 W. Perching represents the robot holding position without movement, consuming only 3.15 W, nearly zero compared to continuous motion. Crawling on a horizontal branch required 13.35 W, while peak power under maximum load reached 34.2 W. The COT during horizontal crawling is 23.4, increasing to 64.6 at a 67.5° incline, reflecting the higher energetic cost of climbing steeper surfaces.
@@
While aerial robots require continuous thrust, the crawler consumes only 3.15 W when perched and 13.35 W during horizontal crawling. A direct comparison with other tree crawlers is left for future work.
This work presented a tether-deployable crawler designed for adaptive locomotion and manipulation within tree canopies. The system combines compliant microspine-based tracks, a dual-track rotary grasper, and an elastic tail, enabling secure attachment and traversal on branches of varying diameters and inclinations. Experiments demonstrated stable gripping on rolls up to 90° and climbing on branches up to 67.5° with a 33% success rate. On horizontal branches, the robot achieved an average speed of 8.33 cm/s (0.55 body lengths/s), with yaw steering up to 10° enabled by the compliant microspine system. Power measurements indicated low energy requirements across modes, with Cost of Transport ranging from 23.4 on horizontal branches to 64.6 at 67.5°, showing efficient operation even on challenging terrain.
Key limitations remain. The current passive wheeled tail provides stability benefits but relies on frictional interaction rather than active force control; biological studies suggest tails enhance balance and stability in arboreal locomotion by modulating angular momentum on narrow supports, which motivates future active designs. Yaw steering is limited to low angles due to reduced contact at higher misalignment, and claims of full aerial deployment are conceptual rather than experimentally demonstrated in this work.
Future work will focus on enhancing tail-assisted stability through force‑ or geometry‑aware control of the gripping angle and active tail actuation to improve robustness on steeper inclinations. Additionally, integrating lightweight sensing and sampling tools will extend autonomous operation within forest canopies, enabling full drone–crawler mission execution in real-world field deployments.